OpenStudy (anonymous):

Form a sequence that has two arithmetic means between -1 and 59. HELP PLEASE!!

OpenStudy (dumbcow):

-1, a, b, 59 and you know that : a-(-1) = b-a = 59-b or using explicit formula \[A_n = A_1 + (n-1)d\] A_1 = -1 A_4 = 59 solve for d

OpenStudy (anonymous):